OpenStudy (anonymous):

What is the square root of 7 over fourth root of 7?

OpenStudy (anonymous):

\[\sqrt{7}/\sqrt[4]{7}\]

OpenStudy (anonymous):

\[7^\frac{ 1 }{4 }\]

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf \frac{\sqrt{7} }{\sqrt[4]{7}} }\)

OpenStudy (solomonzelman):

When you say \(\Huge\color{blue}{ \sf \sqrt{7} }\) it is same as saying \(\Huge\color{blue}{ \sf \sqrt[2]{7} }\)

OpenStudy (jdoe0001):

\(\bf \large {\cfrac{1}{\sqrt[{\color{red} m}]{a^{\color{blue} n}}}= \cfrac{1}{a^{\frac{{\color{blue} n}}{{\color{red} m}}}}\implies a^{-\frac{{\color{blue} n}}{{\color{red} m}}} \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}} \\ \quad \\ \quad \\ \cfrac{\sqrt{7}}{\sqrt[4]{7}}\implies \cfrac{\sqrt{7}}{1}\cdot \cfrac{1}{\sqrt[4]{7}}\implies 7^{\frac{1}{2}}\cdot \cfrac{1}{7^{\frac{1}{4}}}\implies 7^{\frac{1}{2}}\cdot 7^{-\frac{1}{4}} }\)

OpenStudy (solomonzelman):

So, \(\Huge\color{blue}{ \sf \frac{\sqrt{7} }{\sqrt[4]{7}} }\) would be same as \(\Huge\color{blue}{ \sf \frac{\sqrt[2]{7} }{\sqrt[4]{7}} =\frac{7^{1/2}}{7^{1/4} }=7^{1/4}}\)

OpenStudy (solomonzelman):

and \(\Huge\color{blue}{ \sf 7^{1/4}= \sqrt[4]{7^1}=\sqrt[4]{7}}\)

OpenStudy (anonymous):

Thank you so much! I had just figured it out like a minute ago but ts good to know that I got it right! thank you all so much! Would anyone be willing to help me with another one?

OpenStudy (solomonzelman):

You welcome :)

OpenStudy (solomonzelman):

I can try but I stink at math.

OpenStudy (anonymous):

ok great thanks! here it is Which exponential function goes through the points (1, 16) and (4, 128)?

OpenStudy (solomonzelman):

the pattern that I see is (x,y) --> (x+1, y × 2)

OpenStudy (anonymous):

how might that be seen as an exponential function that would go through that line? I have some choices if those might help

OpenStudy (solomonzelman):

Post the choices please.

OpenStudy (anonymous):

here ya go and sorry I should've posted them with the question f(x) = 4(4)x f(x) = 8(2)x f(x) = 8(2)-x f(x) = 4(4)-x

OpenStudy (anonymous):

f(x) = 4(4)^x f(x) = 8(2)^x f(x) = 8(2)^-x f(x) = 4(4)^-x

OpenStudy (solomonzelman):

Yeah that's better.

OpenStudy (anonymous):

is it b?

OpenStudy (solomonzelman):

lets look at the first one. f(x) = 4(4)^x 16= 4(4)^1 64= 4(4)^2 256= 4(4)^3 you can see A doesn't work . Lets check choice B. f(x) = 8(2)^x 16 = 8(2)^1 32 = 8(2)^2 64= 8(2)^3 128 = 8(2)^4 yep it works

OpenStudy (anonymous):

ok thank you soooo much!!

OpenStudy (solomonzelman):

You are welcome !